# Identification and Optimal Nonlinear Control of Turbojet Engine Using Koopman Eigenfunction Model

### v2

@@ 4.2 Optimization of the eigenvalues via temporal projection
For the optimal construction from [45], apriori initial guesses of the eigenvalues need to be supplied. The article mentions a possible optimization of these guesses but it assumes gradient-based approach. However, we want to apply constraints and enforce conditions that must be satisfied by eigenvalues. The penalization terms may not have analytical gradients or their evaluation might be expensive.
The goal is to decompose the estimated autonomous dynamics $f(N)$ to a linear combination of Koopman spectral components given as
@@
or, alternatively, to decompose the trajectories generated by $f(N)$, as
@@
Eq. (33) represents the spatiotemporal decomposition of the dynamics. Note that due to the properties of the linear combination, the eigenfunctions can be scaled so that the initial conditions are $\varphi_{i}(N_{0})=1$. Thus, the temporal decomposition can be rewritten as
@@
The KPDE offers a powerful tool for discovery of the Koopman intrinsic coordinates - eigenfunctions, and thus no trial-and-error procedure connected to DMD-based methods is needed. The problem to be addressed is finding an optimal set of eigenvalues providing a good basis for projection of the autonomous dynamics. This set is composed of the most prominent eigenvalues, most likely with large negative real parts close to 0, as these are the longest persisting in the dynamics. Heavily damped eigenvalues, that will decay almost immediately, are not a suitable choice. A possible option is to identify a large set of eigenfunctions and subsequently utilize regularized LS to find a sparse vector of eigenmodes. For this, the algebraic structure of eigenvalues and eigenfunctions can be effectively exploited. However, it can be a time-consuming process that can often lead to many spurious eigen-tuples [31].
In this paper, a meta-heuristic global optimization method called particle swarm optimization (PSO) was employed to solve the eigenvalue optimization problem since it would be difficult to address the following constraints with gradient-based approaches. This algorithm is a popular choice for global optimization and is supported by the MATLAB Global Optimization Toolbox. For a detailed description, the reader is refer to [52, 53].
The single-spool GTE is a 1st-order system. Thus, the problem can be decoupled to a temporal projection of autonomous trajectories onto trajectories generated by a given set of eigenvalues, such that the projection error is minimized, and subsequent discovery of spatial shapes of eigenfunctions. The former should be reached by optimization of the given set of eigenvalues. Trajectories starting from $k$ different initial conditions are augmented to a single matrix $\mathbf{N_{traj}}=\begin{bmatrix}N_{1}&N_{2}&\cdots&N_{k}\end{bmatrix}^{\textbf{T}}$. The set of Koopman modes should be optimal across the entire operating region. The complete constrained problem is formulated as
@@
$\textbf{E}(\mathbf{\Lambda})=\begin{bmatrix}e^{\mathbf{\lambda}_{1}t}&e^{\mathbf{\lambda}_{2}t}&\cdots&e^{\mathbf{\lambda}_{n}t}\end{bmatrix}^{\textbf{T}},$
$\textbf{K}=\left(\textbf{E}\textbf{E}^{\textbf{T}}-\alpha_{K}\textbf{I}\right)^{-1}\mathbf{EN_{traj}^{T}}\hskip 5.69054pt\bigg{|}_{\mathbf{\Lambda}=\text{const.}},$
@@
where $\alpha_{K}$ is a regularization weight factor.
The objective function addresses the above-mentioned constraints through penalization. It also includes an innovative additional feature that allows to switch between distinct and repeated real eigenvalues, for which the corresponding generalized eigenfunctions obey the following equations:
$\frac{d}{dt}\begin{bmatrix}\varphi_{R}\\ \varphi_{D}\end{bmatrix}=\begin{bmatrix}\lambda&1\\ 0&\lambda\end{bmatrix}\begin{bmatrix}\varphi_{R}\\ \varphi_{D}\end{bmatrix},$ (38)
where $\varphi_{R}$ and $\varphi_{D}$ denote eigenfunctions corresponding to the distinct and repeated eigenvalue, respectively.
@@
The $te^{\lambda_{1}t}$ term is called secular term and corresponds to a transient energy growth. In the context of fitting the autonomous dynamics, secular terms can help adapt the model better to the shape of trajectories.
Generation of purely exponential trajectories (distinct real) or secular terms (repeated eigenvalues) is based on a condition of a minimum distance between two consecutive eigenvalues. Eigenvalues are first sorted in a descending order and distances between each subsequent pair are computed as $\mathbf{\Lambda}(2:\text{end})-\mathbf{\Lambda}(1:\text{end}-1)$. If the distance between a given pair of eigenvalues falls below a specified threshold, the old pair is replaced by a new pair of repeated eigenvalues corresponding to the mean of the old pair.
@@
Regarding the complex eigenvalues, even though the single-spool engine is a strictly dissipative system (RPM monotonously decreases to 0), the slightly oscillatory modes can help to better explain the underlying dynamics as the model can mathematically better conform to the shape of autonomous trajectories. Note that the complex eigenvalue is written as $\lambda=\alpha\pm i\beta$. Thus, we use the notation $\text{Re}(\lambda)=\alpha$ and $\text{Im}(\lambda)=\beta$.
To avoid excessively oscillatory modes, the imaginary-to-real-part ratio $\beta/\alpha$ is checked, and the unsatisfactory modes are penalized according to some function $f(\beta/\alpha)$. The higher the imaginary part, the higher the natural frequency of the mode. Thus, if the ratio is too high, the real part, corresponding to damping, is not sufficiently decaying the amplitude of oscillations. This is summarized in Algorithm 2. In this paper, the function $f(\beta/\alpha)$ is a simple comparison to a specified maximum ratio $(\beta/\alpha)_{max}$, exceeding of which leads to exclusion of the mode.
Furthermore, the principle of switching to repeated complex eigenvalues can also be adopted with the distance now being computed in the complex plane. For this purpose, an upper-triangular distance matrix can be computed, and whenever there is a pair of too close eigenvalues, the secular terms are generated as
$\begin{bmatrix}\text{Re}(\varphi_{R})\\ \text{Im}(\varphi_{R})\\ \text{Re}(\varphi_{D})\\ \text{Im}(\varphi_{D})\end{bmatrix}=e^{\bar{\alpha}t}\begin{bmatrix}\cos{\bar{\beta}}&-\sin{\bar{\beta}}&t\cos{\bar{\beta}}&-t\sin{\bar{\beta}}\\ \sin{\bar{\beta}}&\cos{\bar{\beta}}&t\sin{\bar{\beta}}&t\cos{\bar{\beta}}\\ 0&0&\cos{\bar{\beta}}&-\sin{\bar{\beta}}\\ 0&0&\sin{\bar{\beta}}&\cos{\bar{\beta}}\end{bmatrix}\begin{bmatrix}\text{Re}(\varphi_{R0})\\ \text{Im}(\varphi_{R0})\\ \text{Re}(\varphi_{D0})\\ \text{Im}(\varphi_{D0})\end{bmatrix},$ (40)
where $\bar{\alpha}=(\alpha_{i}+\alpha_{j})/2$ and $\bar{\beta}=(\beta_{i}+\beta_{j})/2$, the initial conditions are all equal to 1, and the distance matrix is given as
@@
     if $\textbf{dist}(\lambda_{i},\lambda_{i+1})$ $<$ tol. then
         $\lambda_{i}=\lambda_{i+1}=(\lambda_{i}+\lambda_{i+1})/2$triangleright Replaces with the mean $\lambda$
@@
         $\textbf{E}_{i+1}=(1+t)e^{\lambda_{i+1}t}$triangleright Secular terms
@@
$\textbf{K}=\left(\textbf{E}\textbf{E}^{\textbf{T}}-\alpha_{K}\textbf{I}\right)^{-1}\textbf{E}\textbf{N}_{\textbf{traj}}^{\textbf{T}}$triangleright Computes projection matrix
$J=\frac{1}{N}\sum_{i=1}^{N}\left(\textbf{N}_{\textbf{traj}}^{\textbf{T}}-\textbf{K}\textbf{E}^{\textbf{T}}\right)$
Algorithm 2 Objective function - complex eigenvalues
@@
$\textbf{E}=\begin{bmatrix}e^{(\alpha\pm i\beta)_{1}t}&e^{(\alpha\pm i\beta)_{2}t}&\cdots&e^{(\alpha\pm i\beta)_{n}t}\end{bmatrix}^{\textbf{T}}$
$\textbf{K}=\left(\textbf{E}\textbf{E}^{\textbf{T}}-\alpha_{K}\textbf{I}\right)^{-1}\textbf{E}\textbf{N}_{\textbf{traj}}^{\textbf{T}}$triangleright Computes projection matrix
$P_{oscil}=f(\beta/\alpha)$triangleright Penalizes excessively oscillatory modes
$J=\frac{1}{N}\sum_{i=1}^{N}\left(\textbf{N}_{\textbf{traj}}^{\textbf{T}}-\textbf{K}\textbf{E}^{\textbf{T}}\right)+P_{oscil}$
@@ 4.3 Identification of analytical eigenfunctions from KPDE
The temporal projection optimization produces optimal eigenvalues and mode amplitudes. Furthermore, by plotting the exponential time trajectories as functions of spool speed, it also reveals the shape of eigenfunctions. However, direct fitting of these shapes was met with problems regarding the accuracy of the final model, especially due to the gradient of eigenfunctions being sensitive to shape discrepancies of the fitted eigenfunctions. Therefore, the KPDE cost function can be minimized instead, and the gradient-based algorithm is utilized for this purpose. The revealed shapes can help with selection of basis functions for the subsequent regression.
@@
In practice, however, $f(N)$ can be arbitrarily complex, especially for highly nonlinear systems such as the GTE. Thus, since the analytical computation can get extremely challenging, it becomes more advantageous to solve Eq. (44) numerically via the method of candidate solutions closely related to the SINDy algorithm.
Instead of utilization of generated autonomous trajectories, the function of autonomous dynamics $f(N)$ is sampled using a nonlinear sampling pattern which secures increasing density of sample points towards 0. Examples of nonlinear sampling approaches are provided in Fig. 2, where the $x$ and $y$ axes represent the relative position of the sample and the corresponding $N$ value, respectively.
@@
Similarly to the $f(N)$ function, the eigenfunctions of a dissipative system should also satisfy $\varphi(0)=0$ for all modes. This is due to the fact, that even the longest persisting mode (largest eigenvalue) should vanish after a finite period of time and thus the $\dot{N}\rightarrow 0$. The nonlinear sampling also indirectly makes the algorithm put more effort to satisfy this condition.
Additionally, as the optimization is two-sided in sense that both the gradient and the eigenfunction (both terms in the error) are optimized simultaneously, it is crucial to prevent the learning to head directly towards the trivial solution $\mathbf{\Xi=0}$. This is avoided by implementation of a penalty that is given by the difference of initial values of eigenfunctions. This way, the cost term acts as an artificial spring, which forces the solution to stay away form the origin of the $\mathbf{\Xi}$ parameter space. The larger the distance, the larger the counterforce. Thus, the complete cost function for a single $\varphi=\mathbf{\Xi\Theta}(N,\textbf{p})$ and distinct real eigenvalue $\lambda$ is given as
$J_{e,r,d}=\left\|\mathbf{\Xi}\frac{\partial\mathbf{\Theta}}{\partial N}\dot{N}-\lambda\mathbf{\Xi\Theta}\right\|_{2}^{2}+\alpha_{1}\|\mathbf{\Xi}|_{1}+\alpha_{2}\|\varphi(N_{0})_{1}-\varphi(N_{0})\|_{2}^{2},$ (45)
where $\alpha_{1}$ and $\alpha_{2}$ are the weights of the L1 sparsity promoting regularization and the trivial solution penalty, respectively, and $\varphi(N_{0})_{1}$ and $\varphi(N_{0})$ represent the initial value of the eigenfunction at the first iteration and at the current iteration, respectively.
@@
For the complex eigenvalues, the KPDE is given as
@@
$Re:\frac{\partial\varphi_{R}}{\partial N}\dot{N}=\alpha\varphi_{R}-\beta\varphi_{I},$ (48)
$Im:\frac{\partial\varphi_{I}}{\partial N}\dot{N}=\beta\varphi_{R}+\alpha\varphi_{I}.$ (49)
@@
$J_{e,c}=\left\|\frac{\partial\varphi_{R}}{\partial N}\dot{N}-\alpha\varphi_{R}+\beta\varphi_{I}\right\|_{F}^{2}+\left\|\frac{\partial\varphi_{I}}{\partial N}\dot{N}-\beta\varphi_{R}-\alpha\varphi_{I}\right\|_{F}^{2},$ (50)

### v6

@@ 4.2 Optimization of the eigenvalues via temporal projection
For the optimal construction from [bib45], a priori initial guesses of the eigenvalues need to be supplied. The article mentions a possible optimization of eigenvalues, but it assumes a gradient-based approach. However, we want to apply constraints and enforce conditions that must be satisfied by eigenvalues. The penalization terms may not have analytical gradients, or their evaluation might be expensive.
The goal is to decompose the estimated autonomous dynamics $f(N)$ into a linear combination of Koopman spectral components given as
@@
alternatively, to decompose the trajectories generated by $f(N)$, as
@@
Equation (35) represents the spatiotemporal decomposition of the dynamics. Note that due to the properties of the linear combination, the eigenfunctions can be scaled so that the initial conditions are $\varphi_{i}(N_{0})=1$. Thus, the temporal decomposition can be rewritten as
@@
The KPDE offers a powerful tool for the discovery of the Koopman intrinsic coordinates - eigenfunctions, and thus, no trial-and-error procedure connected to DMD-based methods is needed. The problem to be addressed is finding an optimal set of eigenvalues providing a good basis for the projection of the autonomous dynamics. This set is composed of the most prominent eigenvalues with negative real parts close to the imaginary axis, as these are the longest persisting in the dynamics. Heavily damped eigenvalues, which will decay almost immediately, are not a suitable choice. A possible option is to identify a large set of eigenfunctions and subsequently utilize regularized LS to find a sparse vector of eigenmodes. For this, the algebraic structure of eigenvalues and eigenfunctions can be effectively exploited. However, it can be a time-consuming process that can often lead to many spurious eigen-tuples [bib31].
In this paper, a meta-heuristic global optimization method called particle swarm optimization (PSO) was employed to solve the eigenvalue optimization problem since it would be difficult to address the following constraints with gradient-based approaches. This algorithm is a popular choice for global optimization and is supported by the MATLAB Global Optimization Toolbox. For a detailed description, the reader is referred to [bib52, bib53].
The single-spool GTE is a 1st-order system. Thus, the problem can be decoupled into a temporal projection of autonomous trajectories onto trajectories generated by a given set of eigenvalues, such that the projection error is minimized, and subsequent discovery of spatial shapes of eigenfunctions. The former should be reached by optimization of the given set of eigenvalues. Trajectories starting from $k$ different initial conditions are augmented to a single matrix $\mathbf{N_{traj}}=\begin{bmatrix}N_{1}&N_{2}&\cdots&N_{k}\end{bmatrix}^{\intercal}$. The set of Koopman modes should be optimal across the entire operating region. The complete constrained problem is formulated as
@@
$\textbf{E}(\mathbf{\Lambda})=\begin{bmatrix}e^{\mathbf{\lambda}_{1}t}&e^{\mathbf{\lambda}_{2}t}&\cdots&e^{\mathbf{\lambda}_{n}t}\end{bmatrix}^{\intercal}$,
$\textbf{K}^{\intercal}=\left(\textbf{E}\textbf{E}^{\intercal}+\alpha_{K}\textbf{I}\right)^{-1}\mathbf{EN_{traj}}^{\intercal}\hskip 5.69054pt\bigg|_{\mathbf{\Lambda}=\text{const.}},$
@@
where K is the matrix of Koopman modes associated with the given set of eigenvalues, $\mathbf{\Lambda}$, $\alpha_{K}$ is a regularization weight factor, and $\text{Re}(\mathbf{\Lambda})_{\text{max}}$ ensures a sufficient stability margin.
The objective function addresses the above-mentioned constraints through penalization. It also includes an innovative additional feature that allows for switching between distinct and repeated real eigenvalues, for which the corresponding generalized eigenfunctions obey the following equations:
$\frac{\text{d}}{\text{d}t}\begin{bmatrix}\varphi_{R}\\ \varphi_{D}\end{bmatrix}=\begin{bmatrix}\lambda&1\\ 0&\lambda\end{bmatrix}\begin{bmatrix}\varphi_{R}\\ \varphi_{D}\end{bmatrix}$, (40)
where $\varphi_{R}$ and $\varphi_{D}$ denote eigenfunctions corresponding to the distinct and repeated eigenvalues, respectively.
@@
The $te^{\lambda_{1}t}$ term is called the secular term and corresponds to a transient energy growth. In the context of fitting the autonomous dynamics, secular terms can help adapt the model better to the shape of trajectories.
Generation of purely exponential trajectories (distinct real) or secular terms (repeated eigenvalues) is based on a condition of a minimum distance between two consecutive eigenvalues. Eigenvalues are first sorted in descending order and distances between each subsequent pair are computed as $\mathbf{\Lambda}(2:\text{end})-\mathbf{\Lambda}(1:\text{end}-1)$. If the distance between a given pair of eigenvalues falls below a specified threshold, the old pair is replaced by a new pair of repeated eigenvalues corresponding to the mean of the old pair.
@@
Regarding the complex eigenvalues, even though the single-spool engine is a strictly dissipative system (RPM monotonously decreases to 0), the slightly oscillatory modes can help to better explain the underlying dynamics, as the model can mathematically better conform to the shape of autonomous trajectories. Note that the complex eigenvalue is written as $\lambda=\alpha\pm i\beta$. Thus, we use the notation $\text{Re}(\lambda)=\alpha$ and $\text{Im}(\lambda)=\beta$.
To avoid excessively oscillatory modes, the imaginary-to-real-part ratio $\beta/\alpha$ is checked, and the unsatisfactory modes are penalized according to some function $f(\beta/\alpha)$. The higher the imaginary part, the higher the natural frequency of the mode. Thus, if the ratio is too high, the real part, corresponding to damping, does not sufficiently decay the amplitude of oscillations. This is summarized in Algorithm 2. In this paper, the function is
$f(\beta/\alpha)=(\beta/\alpha)_{max},$ (42)
exceeding which leads to exclusion of the mode.
Furthermore, the principle of switching to repeated complex eigenvalues can also be adopted, with the distance now being computed in the complex plane. For this purpose, an upper-triangular distance matrix D can be computed, and whenever there is a pair of too close eigenvalues, the secular terms are generated as
$\begin{bmatrix}\Re(\varphi_{R})\\ \Im(\varphi_{R})\\ \Re(\varphi_{D})\\ \Im(\varphi_{D})\end{bmatrix}=e^{\bar{\alpha}t}\begin{bmatrix}C({\bar{\beta}})&-S({\bar{\beta}})&tC({\bar{\beta}})&-tS({\bar{\beta}})\\ S({\bar{\beta}})&C({\bar{\beta}})&tS({\bar{\beta}})&tC({\bar{\beta}})\\ 0&0&C({\bar{\beta}})&-S({\bar{\beta}})\\ 0&0&S({\bar{\beta}})&C({\bar{\beta}})\end{bmatrix}$, (43)
where $\bar{\alpha}=(\alpha_{i}+\alpha_{j})/2$ and $\bar{\beta}=(\beta_{i}+\beta_{j})/2$, $S({\bar{\beta}})$ and $C({\bar{\beta}})$ denote $\sin\bar{\beta t}$ and $\cos\bar{\beta t}$, respectively. The initial conditions are all equal to 1, and the distance matrix is given as
@@
  if $\textbf{dist}(\lambda_{i},\lambda_{i+1})$ << tol. then
   $\lambda_{i}=\lambda_{i+1}=(\lambda_{i}+\lambda_{i+1})/2$ $\triangleright$ Replaces with the mean $\lambda$
@@
   $\textbf{E}_{i+1}=(1+t)e^{\lambda_{i+1}t}$ $\triangleright$ Secular terms
@@
$\textbf{K}^{\intercal}=\left(\textbf{E}\textbf{E}^{\intercal}+\alpha_{K}\textbf{I}\right)^{-1}\textbf{E}\textbf{N}_{\textbf{traj}}^{\intercal}$ $\triangleright$ Computes projection matrix
$J=\left\|\textbf{N}_{\textbf{traj}}-\textbf{K}\textbf{E}\right\|_{F}^{2}$
Algorithm 2 Objective function - distinct complex eigenvalues
@@
$\textbf{E}=\begin{bmatrix}e^{(\alpha\pm i\beta)_{1}t}&e^{(\alpha\pm i\beta)_{2}t}&\cdots&e^{(\alpha\pm i\beta)_{n}t}\end{bmatrix}^{\intercal}$
$\textbf{K}^{\intercal}=\left(\mathbf{EE^{\intercal}}+\alpha_{K}\textbf{I}\right)^{-1}\textbf{E}\textbf{N}_{\textbf{traj}}^{\intercal}$ $\triangleright$ Computes projection matrix
$P_{oscil}=f(\beta/\alpha)$ $\triangleright$ Penalizes excessively oscillatory modes
$J=\left\|\textbf{N}_{\textbf{traj}}-\textbf{K}\textbf{E}\right\|_{F}^{2}+P_{oscil}$
@@ 4.3 Identification of analytical eigenfunctions from KPDE
The temporal projection optimization produces optimal eigenvalues and mode amplitudes. Furthermore, by plotting the exponential time trajectories as functions of spool speed, it also reveals the shape of eigenfunctions. However, direct fitting of these shapes was met with problems regarding the accuracy of the final model, especially due to the gradient of eigenfunctions being sensitive to shape discrepancies of the fitted eigenfunctions. Therefore, the KPDE cost function can be minimized instead, and the gradient-based algorithm is utilized for this purpose. The revealed shapes can help with the selection of basis functions for the subsequent regression.
@@
In practice, however, $f(N)$ can be arbitrarily complex, especially for highly nonlinear systems such as the GTE. Thus, since the analytical computation can get extremely challenging, it becomes more advantageous to solve (46) numerically via the method of candidate solutions closely related to the SINDy algorithm.
Instead of the utilization of generated autonomous trajectories, the function of autonomous dynamics $f(N)$ is sampled using a nonlinear sampling pattern that ensures an increasing density of sample points towards 0. Examples of nonlinear sampling approaches are provided in Fig. 3, where the horizontal and vertical axes represent the relative position of the sample and the corresponding $N$ value, respectively.
@@
Similarly to the $f(N)$ function, the eigenfunctions of a dissipative system should also satisfy $\varphi(0)=0$ for all modes. This is because even the longest persisting mode (largest eigenvalue) should vanish after a finite time and thus the $\dot{N}\rightarrow 0$. The nonlinear sampling also indirectly makes the algorithm put more effort into satisfying this condition.
Additionally, as the optimization is two-sided in the sense that both the gradient and the eigenfunction (both terms in the error) are optimized simultaneously, it is crucial to prevent the learning from heading directly towards the trivial solution $\mathbf{\Xi=0}$. This is avoided by a penalty term that is a function of the difference between the value of the eigenfunction at $N_{0}$ in the current iteration and the value at $N_{0}$ in the first iteration. This way, the cost term acts as an artificial spring, which forces the solution to stay away from the origin of the $\mathbf{\Xi}$ parameter space. The larger the distance, the larger the counterforce. Thus, the complete cost function for a single $\varphi=\mathbf{\Xi\Theta}(N,\textbf{p})$ and distinct real eigenvalue $\lambda$ is given as
$J_{e,r,d}=$ $\left\|\mathbf{\Xi}\frac{\partial\mathbf{\Theta}}{\partial N}\dot{N}-\lambda\mathbf{\Xi\Theta}\right\|_{2}^{2}+\alpha_{1}\|\mathbf{\Xi}\|_{1}$
$+\alpha_{2}\|\varphi(N_{0})_{1}-\varphi(N_{0})\|_{2}^{2},$ (48)
where $\alpha_{1}$ and $\alpha_{2}$ are the weights of the L1 sparsity-promoting regularization and the trivial solution penalty, respectively, and $\varphi(N_{0})_{1}$ and $\varphi(N_{0})$ represent the initial value of the eigenfunction at the first iteration and the current iteration, respectively.
@@
For the complex eigenfunctions, the KPDE is given as
@@
$\text{Re}:\frac{\partial\varphi_{R}}{\partial N}\dot{N}=\alpha\varphi_{R}-\beta\varphi_{I},$ (51)
$\text{Im}:\frac{\partial\varphi_{I}}{\partial N}\dot{N}=\beta\varphi_{R}+\alpha\varphi_{I}.$ (52)
@@
$J_{e,c}=$ $\left\|\frac{\partial\varphi_{R}}{\partial N}\dot{N}-\alpha\varphi_{R}+\beta\varphi_{I}\right\|_{F}^{2}$
$+\left\|\frac{\partial\varphi_{I}}{\partial N}\dot{N}-\beta\varphi_{R}-\alpha\varphi_{I}\right\|_{F}^{2},$ (53)